OpenStudy (melodycrazysouthern):

Complete the definition of an irrational number below. An irrational number is a -- -- decimal. Unlike rational numbers, irrational numbers -- -- be expressed as a fraction,p/q , where p and q are -- --.

OpenStudy (mathmale):

I'd strongly suggest that you look up "irrational number" as well as "rational number." Look for examples. You'll need to know and understand this definition, and so the time spent looking it up would be worthwhile.

OpenStudy (melodycrazysouthern):

ive tried still dont get it

OpenStudy (rcp031):

Rational number= 10/3=3.33333 repeating decimals While Irrational number=22/7=3.14285714....... Non repeating decimals

OpenStudy (rcp031):

p & q are simply any integers

OpenStudy (sapphiremoon):

Non-terminating and non-repeating, like @rcp031 said. Numbers like pi and e are irrational because they don't stop and don't continuously repeat (and they can't be represented as fractions).

OpenStudy (melodycrazysouthern):

thanks guys

OpenStudy (mathmale):

Actually, a rational number is one that can be expressed as the ratio of two integers: 9/10, 4/5, 1/3, and so on. The quantity pi cannot be so expressed, and thus is irrational.

OpenStudy (sapphiremoon):

@mathmale Isn't that what I said?

OpenStudy (melodycrazysouthern):

plz dont argue

OpenStudy (sapphiremoon):

@melodycrazysouthern Were we arguing? I wasn't aware... Just asking a question from my point of view...